# ChamaleonLLM: Batch-Aware Dynamic Low-Rank Adaptation via Inference-Time Clusters

### v1

@@ 1 Introduction
ChamaleonLLM is designed to address this limitation by enabling the model to adapt its decoder weights during inference based on the structure and statistics of the input batch. Inputs are grouped into clusters based on their semantic and syntactic similarities. By leveraging precomputed token embeddings, the inference engine identifies coherent groups within each batch, ensuring that similar examples are processed together. Rather than relying on pre-learned and fixed uniforms (changeable masks) or static LoRA updates, ChamaleonLLM employs a hyper-network to generate low-rank adaptation parameters in real-time. This dynamic generation is based on the aggregated statistics of the clustered batch, allowing the model to tailor its adaptations to the prevailing context. This approach not only removes the necessity for multiple expert models or a vast wardrobe of pre-stored masks but also leverages the collective context of the batch, leading to enhanced performance across a range of tasks. The primary contributions of this work are three-fold: (1) We introduce a method to dynamically generate low-rank updates based on batch statistics, resulting in a self-adaptive model during inference. (2) We employ a clustering-based on normalized token embeddings that groups similar inputs, ensuring context-aware adaptation. (3) We provide experimental results demonstrating the superiority of ChamaleonLLM over traditional LoRA fine-tuning.
@@ 2 Related Work
LoRA has emerged as a popular method for fine-tuning large pre-trained models with minimal computational resources. The fundamental idea behind LoRA is to freeze the pre-trained weights and introduce a pair of trainable, low-rank matrices that approximate the necessary updates for a given task. By operating in a low-dimensional subspace, LoRA significantly reduces the number of trainable parameters while enabling effective adaptation. Despite its success, standard LoRA is inherently static during inference. Once the low-rank updates are learned, they are fixed and applied uniformly, regardless of the heterogeneity of the incoming data. This static nature can limit the model’s responsiveness to dynamic input distribution or context changes. An alternative to direct low-rank adaptation is using changeable masks as pre-learned uniforms (Sun et al., 2025). This method involve training a set of masks—each tailored to specific tasks or data distributions—that can be blended or selected during inference. The blending is implemented through a softmax-based selector mechanism that chooses the most appropriate mask based on a task embedding or a similar indicator. While this method introduces a degree of flexibility, it comes with drawbacks:
@@
  • Static Adaptation: Even though blending introduces some variability, the masks are fine-tuned individually and not co-optimized for joint adaptation. This can lead to suboptimal performance when the masks are applied in a blended fashion.
@@
Similarly, another recent approach Tan et al. (2024) introduces a hypernetwork‐based method for dynamic layer operations, which generates context‑dependent low‑rank updates by adjusting the model’s depth on a per‐sample basis; where the hypernetwork adapts the network structure for each individual input—effectively generating adjustments for model depth dynamically at inference time. In contrast to per-sample or depth-focused adaptation methods, our current work—ChamaleonLLM—exploits batch-level context to generate adaptive low-rank updates. Rather than adapting each sample independently, we cluster similar inputs and compute aggregated token embedding statistics across the batch; these statistics then drive our hyper-network to generate low-rank updates that reflect the collective context. ChamaleonLLM achieves a coherent and context-aware adaptation that naturally overcomes the limitations of fixed, per-sample updates and the storage overhead associated with large sets of pre-learned masks, such as the impact of noisy or outlier samples, and the storage overhead associated with maintaining pre-learned masks.
@@ 3 Methodology
ChamaleonLLM is built upon a pre-trained causal language model (e.g., GPT‑2), where the conventional transformer layers and language modeling head are augmented with low-rank adaptation modules. The innovation lies in the dynamic generation of low-rank parameters based on batch-aware statistics. This process involves two major components: (1) batch-aware clustering and (2) adaptive low-rank update generation. The central hypothesis is that inputs arriving in a batch often exhibit significant similarity—whether semantic, syntactic, or stylistic. By grouping similar inputs, we can extract shared context and exploit common features in the batch. This grouping not only aids in noise reduction by averaging but also ensures that the low-rank updates are tailored to the collective characteristics of the batch rather than isolated samples. Using the normalized token embeddings, we perform k‑means clustering. The number of clusters is chosen based on the batch size to ensure each cluster is large enough to provide robust statistics while capturing meaningful variations between groups. The clustering algorithm minimizes the Euclidean distance between points and cluster centroids, and through iterative refinement, it assigns each input to a cluster that best represents its features. Once the clusters are defined, the batch is reconstructed such that each mini-batch contains inputs from the same cluster. This reorganization is critical because it ensures that subsequent low-rank adaptations are computed on a homogenous set of inputs.
@@
The core idea of dynamic adaptation is to generate low-rank modifications specific to the current batch’s context. This work proposes a more flexible alternative to LoRA that uses a hyper-network. Instead of applying a fixed set of low-rank updates (as in standard LoRA), ChamaleonLLM employs a hyper-network to produce these parameters on-the-fly. The hyper-network takes as input the mean of the token embeddings from the entire batch, processes this aggregated statistic through several fully connected layers with non-linear activations, and outputs the parameters for the low-rank update. This on-the-fly generation allows the model to dynamically adjust to the specifics of the batch, capturing nuances that a static update might miss. The overall architecture maintains the integrity of the pre-trained model while introducing minimal additional parameters. The transformer layers are wrapped with LoRA modules that operate similarly to conventional implementations, ensuring that most of the model’s capacity remains unchanged. For the LM head, the hyper-network variant ensures that adaptation is context-dependent. During inference, the model first processes the batch to compute the necessary embeddings to cluster the inputs and then generates a custom low-rank update for the LM head based on the cluster’s aggregated statistics. The final output is produced by applying this adapted LM head to the last hidden states of the transformer.
@@ 4 Experiments
For our experiments, we use the WikiText-2 raw dataset, which provides a challenging benchmark due to its diverse and natural language texts. The dataset is split into training and validation sets. Each text sample is tokenized using a pre-trained tokenizer (of the GPT‑2). Tokenization includes truncation to a maximum length (256 tokens) and padding to ensure uniform sequence lengths. LM input embeddings (token + positional) and token embeddings (for clustering) are computed for every example. The token embeddings are normalized to facilitate robust clustering. A key part of our methodology is creating data loaders that reflect the clustered nature of the inputs. Using the precomputed normalized token embeddings, we apply k‑means clustering. The number of clusters is derived based on the desired batch size and overall dataset size. After clustering, indices are grouped so that each mini-batch contains examples from a single cluster. This ensures that each batch is contextually coherent. We implement a custom data loader that leverages these clusters as batch samplers. The data loader returns batches fed into the model during training and evaluation.
@@
Unadapted GPT-2 124,439,808 10.2921 10.2513 28,319
Traditional LoRA 204,100 1.4088 1.3528 3.8683
@@
  • ChamaleonLLM Fine-Tuning: In this regime, while the transformer layers continue to use static LoRA modules, the LM head is adapted by a hyper-network LoRA module that generates low-rank update parameters based on the mean token embeddings of the batch (Ha et al., 2016). The loss is computed on the LM head outputs, and optimized similarly.
The evaluation is performed at the end of each epoch on the validation set, where batches are processed through the clustering pipeline, and the average loss is computed. This systematic evaluation ensures that our comparisons are fair and reflect real-world performance. Experimental results demonstrate that we achieved significant improvements over the traditional LoRA method. ChamaleonLLM consistently achieves a lower average validation loss than the default LoRA approach. This suggests that dynamic adaptation based on batch statistics leads to better generalization. All codes for ChamaleonLLM are open-sourced to ensure the reproducibility of experimental results.
@@ 5 Conclusion
In this paper, we presented ChamaleonLLM—a novel framework that enables inference-time adaptation of large language models through batch-aware clustering and dynamic low-rank parameter generation. By clustering similar inputs and using a hyper-network to generate low-rank updates based on the aggregated batch statistics, our method provides a flexible and efficient alternative to static fine-tuning approaches such as traditional LoRA or uniform mask blending. Our extensive experiments on WikiText-2 demonstrate that ChamaleonLLM achieves lower validation loss and perplexity than baseline methods with dynamic batch-aware contextual adaptation. The proposed framework reduces the need for storing multiple expert masks and adapts to diverse input distributions in real-time, making it highly suitable for modern, dynamic inference environments. Ultimately, ChamaleonLLM represents a promising step toward truly self-adaptive language models capable of adjusting to the ever-changing demands of natural language processing.
The success of ChamaleonLLM can largely be attributed to its exploitation of batch-level context. By averaging token embeddings over a batch, the hyper-network input becomes less sensitive to individual outliers, leading to more robust parameter generation. Clustering groups together semantically or stylistically similar inputs allows the hyper-network to extract stronger, more relevant signals for adaptation. Although the hyper-network introduces additional computational steps during inference, the overhead is modest compared to the overall inference time. The trade-off between increased computation and improved performance is highly favorable. Our approach is not tied to pre-defined task categories. Instead, it adapts to the nuances in the data at inference time, offering a versatile solution for open-domain tasks. Although ChamaleonLLM is designed for batch-level context adaptation, integrating instruction-based or task-specific signals could further enhance performance on specialized tasks. Additional experiments on diverse sets of datasets and tasks will be necessary to fully evaluate the scalability and robustness of our method in production environments.

### v2

@@ 1 Introduction
ChamaleonLLM is designed to address this limitation by enabling the model to adapt its decoder weights during inference based on the structure and statistics of the input batch. Inputs are grouped into clusters based on their semantic and syntactic similarities. By leveraging precomputed token embeddings, the inference engine identifies coherent groups within each batch, ensuring that similar examples are processed together. Rather than relying on pre-learned and fixed uniforms (changeable masks) or static LoRA updates, ChamaleonLLM employs a hyper-network (Ha et al., 2016) to generate low-rank adaptation parameters in real-time. This dynamic generation is based on the aggregated statistics of the clustered batch, allowing the model to tailor its adaptations to the prevailing context. This approach not only removes the necessity for multiple expert models or a vast wardrobe of pre-stored masks but also leverages the collective context of the batch, leading to enhanced performance across a range of tasks. The primary contributions of this work are three-fold: (1) We introduce a method to dynamically generate low-rank updates based on batch statistics, resulting in a self-adaptive model during inference. (2) We employ a clustering-based on normalized token embeddings that groups similar inputs, ensuring context-aware adaptation. (3) We provide experimental results demonstrating the superiority of ChamaleonLLM over traditional LoRA fine-tuning.
@@ 2 Related Work
LoRA has emerged as a popular method for fine-tuning large pre-trained models with minimal computational resources. The fundamental idea behind LoRA is to freeze the pre-trained weights and introduce a pair of trainable, low-rank matrices that approximate the necessary updates for a given task. By operating in a low-dimensional subspace, LoRA significantly reduces the number of trainable parameters while enabling effective adaptation. Despite its success, standard LoRA is inherently static during inference. Once the low-rank updates are learned, they are fixed and applied uniformly, regardless of the heterogeneity of the incoming data. This static nature can limit the model’s responsiveness to dynamic input distribution or context changes. Several methods have been proposed that explore dynamic and parameter-efficient adaptation from various perspectives. For instance, Chen et al. (2024) introduced a multiplicative sparse factorization technique that leverages the low-rank structure in fine-tuning, showing that a single adjustment may not be sufficient for robust performance. Similarly, Kim et al. (2024) demonstrated that a single linear layer can yield task-adapted low-rank matrices, highlighting the potential for dynamic low-rank adaptations. Xiao et al. (2023) with task-agnostic low-rank adapters for unseen English dialects, has also demonstrated the power of adaptive methods by addressing robustness across diverse linguistic scenarios.
Building on these advancements, recent work has sought to further enhance adaptability through more modular and context-sensitive designs. Ostapenko et al. (2024) proposed a modular approach for LLMs by building and reusing a library of LoRAs. Unlike standard LoRA—which applies a single static low-rank update during inference—this modular method dynamically selects or combines specialized LoRA modules based on the task or context, thereby enabling a more flexible and responsive adaptation to diverse and evolving input distributions. In contrast, an alternative strategy employs pre-learned changeable masks (Sun et al., 2025). This approach involves training a set of masks—each tailored to specific tasks or data distributions—that are blended or selected during inference via a softmax-based selector informed by a task embedding or similar indicator. While these methods introduce dynamic adjustment, they come with some practical limitations, such as:
@@
  • Adaptation Flexibility: Even though blending introduces some variability, the masks are fine-tuned individually and not co-optimized for joint adaptation. This can lead to suboptimal performance when the masks are applied in a blended fashion.
@@
Another line of research has focused on hypernetwork-based task conditioning and multi-task learning. As an early attempt, Mahabadi et al. (2021) proposed a shared hypernetwork framework for parameter‑efficient multi‑task fine‑tuning. Later, He et al. (2022) introduced Hyperprompt, a method that conditions transformers using hypernetwork‑generated prompts, while Ivison & Peters (2022) presented hyperdecoders capable of producing instance‑specific decoders for multi‑task NLP. More recently, Tan et al. (2024) proposed a strategy that dynamically adjusts model depth on a per‑sample basis by generating context‑dependent low‑rank updates via a hypernetwork, and Schug et al. (2024) reinterpreted the attention mechanism as a hypernetwork, embedding dynamic parameter generation in the model’s core architecture. Lastly, Ortiz-Barajas et al. (2024) integrated hypernetwork‑based LoRA and adapter layers into multi‑task transformers for sequence labeling, while Lv et al. (2024) leveraged constrained low‑rank adapter generation to achieve efficient cross‑task generalization.
@@
To sum up, in contrast to per-sample or depth-focused adaptation methods, ChamaleonLLM exploits batch-level context to generate adaptive low-rank updates. Rather than adapting each sample independently, we cluster similar inputs and compute aggregated token embedding statistics across the batch; these statistics then drive our hyper-network to generate low-rank updates that reflect the collective context. ChamaleonLLM achieves a coherent and context-aware adaptation that naturally overcomes the limitations of fixed, per-sample updates and the storage overhead associated with large sets of pre-learned masks, such as the impact of noisy or outlier samples, and the storage overhead associated with maintaining pre-learned masks. To conclude, while traditional methods like static LoRA and changeable masks offer efficient adaptation strategies, the recent surge in dynamic approaches demonstrates a growing consensus: incorporating context—whether at the sample or task level—can significantly enhance the adaptability and robustness of LLMs. Our work builds on previous research by uniquely leveraging batch-level clustering to generate context-aware low-rank updates, thereby improving the hypernetwork-based adaptation using aggregated batch statistics.
@@ 3 Methodology
In this work, we introduce ChamaleonLLM, a dynamic adaptation framework leveraging batch-aware clustering and hyper-network–driven low-rank adaptations to efficiently tailor the model to the input context. ChamaleonLLM can be built upon any pre-trained causal language model (e.g., GPT‑2) by augmenting the conventional transformer layers and language modeling head with low-rank adaptation modules. The innovation lies in the dynamic generation of low-rank parameters based on batch-aware statistics. This process involves two major components: (1) batch-aware clustering and (2) adaptive low-rank update generation. The central hypothesis is that inputs arriving in a batch often exhibit significant similarity—whether semantic, syntactic, or stylistic. By grouping similar inputs, we can extract shared context and exploit common features in the batch. This grouping not only aids in noise reduction by averaging but also ensures that the low-rank updates are tailored to the collective characteristics of the batch rather than isolated samples. Using the normalized token embeddings, we perform k‑means clustering. The number of clusters is chosen based on the batch size to ensure each cluster is large enough to provide robust statistics while capturing meaningful variations between groups. The clustering algorithm minimizes the Euclidean distance between points and cluster centroids, and through iterative refinement, it assigns each input to a cluster that best represents its features. Once the clusters are defined, the batch is reconstructed such that each mini-batch contains inputs from the same cluster. This reorganization is critical because it ensures that subsequent low-rank adaptations are computed on a homogenous set of inputs.
@@
The core idea of dynamic adaptation is to generate low-rank modifications specific to the current batch’s context. This work proposes a more flexible alternative to LoRA that uses a hyper-network. Instead of applying a fixed set of low-rank updates (as in standard LoRA), ChamaleonLLM employs a hyper-network to produce these parameters on-the-fly. The hyper-network takes as input the mean of the token embeddings from the entire batch, processes this aggregated statistic through several fully connected layers with non-linear activations, and outputs the parameters for the low-rank update. This on-the-fly generation allows the model to adjust to the specifics of the batch dynamically, capturing nuances that a static update might miss. The overall architecture maintains the integrity of the pre-trained model while introducing minimal additional parameters. The transformer layers are wrapped with LoRA modules that operate similarly to conventional implementations, ensuring that most of the model’s capacity remains unchanged. For the LM head, the hyper-network variant ensures that adaptation is context-dependent. During inference, the model first processes the batch to compute the necessary embeddings to cluster the inputs and then generates a custom low-rank update for the LM head based on the cluster’s aggregated statistics. The final output is produced by applying this adapted LM head to the last hidden states of the transformer.
3.1 Implementation
To leverage the inherent similarity between samples, we first cluster the normalized token embeddings using the k-means algorithm. This ensures that mini-batches are contextually coherent and provide reliable aggregate statistics for conditioning the hyper-network. The number of clusters is determined dynamically (typically by dividing the dataset size by the batch size) to ensure each cluster is sufficiently large to provide reliable aggregate statistics while preserving the similarity among examples. Once clustering is complete, we construct the training and validation data loaders that generate mini-batches composed of homogeneous examples from the same cluster. This clustered batching strategy is critical because it ensures that the subsequent dynamic adaptation via the hyper-network is conditioned on inputs with similar features, enhancing the adaptability and robustness of the generated low-rank updates. By integrating detailed data preprocessing, batch-aware clustering, and dynamic low-rank adaptation, our methodology efficiently leverages the strengths of a pre-trained transformer backbone while adapting to the specific context of each batch.
In our architecture, the transformer layers are augmented with the standard LoRA modules, while the LM head is replaced with a hyper-network LoRA module. Designed specifically for the language modeling (LM) head, this module dynamically predicts update matrices based on the aggregated token embeddings from the current batch. The hyper-network processes the mean token embedding through several fully connected layers with non-linear activations (e.g., ReLU) and dropout, ultimately outputting the low-rank matrices (denoted as A and/or B). Initialization is performed using Xavier and small-scale normal initialization. A dedicated procedure handles wrapping by freezing the original model parameters and applying the appropriate LoRA modifications. This ensures that only the LoRA and hyper-network components are trainable, thereby preserving the representational capacity of the pre-trained model. The model is trained using a standard next-token prediction objective with cross-entropy loss. During each training iteration, the following steps are executed:
  1. 1.
    Transformer Processing: The input batch of clustered examples, is passed through the frozen transformer layers augmented with LoRA, yielding hidden states for each token.
  2. 2.
    Dynamic LM Head Adaptation: The last hidden state is passed to the LM head. Simultaneously, the LM head receives the aggregated token embeddings (computed as the mean over the batch) as input, allowing the hyper-network to generate dynamic low-rank updates.
  3. 3.
    Prediction and Loss Computation: The adapted LM head produces logits for next-token prediction, which are compared against the target tokens to compute the cross-entropy loss.
  4. 4.
    Optimization: Gradients are computed using the AdamW optimizer, and only the LoRA and hyper-network parameters are updated.
@@ 4 Experiments
For our experiments, we have used WikiText-2 (Merity et al., 2016) and Alpaca (Taori et al., 2023) datasets, providing meaningful benchmarks due to their diverse and natural language texts. The datasets are split into training and validation sets. Each text sample is tokenized using a pre-trained GPT-2 tokenizer by truncating to a maximum length and padding to ensure uniform sequence lengths. LM input token embeddings are computed for every example, and normalized to facilitate robust clustering. For the Alpaca dataset, we calculate the token embeddings from the instruction part of the input prompts. A crucial part of ChamaleonLLM is creating data loaders that reflect the clustered nature of the inputs. Using the precomputed token embeddings, we apply k‑means clustering with the number of clusters based on the desired batch size and overall dataset size. After clustering, indices are grouped so each mini-batch contains examples from a single cluster. This ensures that each batch is contextually coherent. We implement a custom data loader that leverages these clusters as batch samplers. The data loader returns batches fed into the model during training and evaluation. The evaluation is performed at the end of each epoch on the validation set, where batches are processed through the clustering pipeline, and the average loss is computed. This systematic evaluation ensures that our comparisons are fair and reflect real-world performance. All codes for ChamaleonLLM are open-sourced to ensure the reproducibility of experimental results.
@@
Unadapted GPT-2 124,439,808 13.8144 13.7876 972,500
Traditional LoRA 204,100 0.5504 0.5023 1.6525
@@
Table 2: Comparison of Low-Rank Adaptation Regimes on Alpaca Dataset
Unadapted GPT-2 124,439,808 13.3881 13.3881 652,166
Traditional LoRA 204,100 0.1254 0.1527 1.1650
ChamaleonLLM* 6,786,596 0.0810 0.1082 1.1143
    Token embeddings for adaptation are calculated from the instruction part of the prompt only.
@@
  • ChamaleonLLM Fine-Tuning: In this regime, while the transformer layers continue to use static LoRA modules, the LM head is adapted by a hyper-network LoRA module that generates low-rank update parameters based on the mean token embeddings of the batch. The loss is computed on the LM head outputs, and optimized similarly.
Experimental results demonstrate that ChamaleonLLM significantly improves over the traditional LoRA approach, consistently achieving a lower average validation loss on both datasets, and suggest that dynamic adaptation based on batch statistics leads to better convergence and generalization. As shown in Table 1, ChamaleonLLM achieves a validation loss reduction of approximately 25% compared to traditional LoRA, along with a perplexity improvement of roughly 12%. Table 2 further indicates that it reduces the validation loss by almost 30% on instruction fine-tuning. These improvements underscore the effectiveness of the batch-level context adaptation strategy and can largely be attributed to the exploitation of batch-level context. By averaging token embeddings over each batch, the hyper-network input becomes less sensitive to individual outliers, leading to robust low-rank parameter generation. Clustering groups together semantically or stylistically similar inputs allows the hyper-network to extract stronger, more relevant signals for adaptation.
Although ChamaleonLLM has more trainable parameters than LoRA, this increase is justified by many practical benefits, as it dynamically generates low-rank updates on-the-fly rather than training and storing separate adaptation matrices for each sample or task. Avoiding the storage of numerous pre-learned matrices significantly reduces memory and processing overheads during inference. The additional parameters enable context-aware, batch-level adaptation, leading to better performance. Despite it also introduces some extra computational steps during inference, the overhead is modest compared to the overall inference time. The trade-off between increased computation and improved performance is highly favorable. Our approach is not tied to pre-defined task categories and adapts to the nuances in the data at inference time, offering a versatile solution for open-domain tasks.
@@ 5 Conclusion
In this paper, we presented ChamaleonLLM—a novel framework that enables inference-time adaptation of large language models through batch-aware clustering and dynamic low-rank parameter generation. By clustering similar inputs and using a hyper-network to generate low-rank updates based on the aggregated batch statistics, our method provides a flexible and efficient alternative to static fine-tuning approaches such as traditional LoRA or uniform mask blending. Our extensive experiments on WikiText-2 and Alpaca demonstrate that ChamaleonLLM achieves lower validation loss and perplexity than baseline methods with dynamic batch-aware contextual adaptation. The proposed framework reduces the need for storing multiple expert masks and adapts to diverse input distributions in real-time, making it highly suitable for modern, dynamic inference environments.